OpenStudy (anonymous):

here all, i need an function for that, in a 3d world, when a weopon fired, the bullet goes with a drift i mean they way it fallows is not a line, so how can we make a function to count how the bullet goes to a target :)

OpenStudy (anonymous):

and it should also count the ricochet ^^

OpenStudy (mathteacher1729):

Sounds like a projectile motion problem. Here is a decent illustration. http://www.staff.amu.edu.pl/~romangoc/M2-1-projectile-motion.html \(v_0\) = initial velocity \(v_0\sin(\theta) \) = Y component of velocity \[v_0\cos(\theta)\] = X component of velocity Gravity is 9.8 m/s^2 in the downward direction. Hope this helps. :)

OpenStudy (anonymous):

thanks ill abstract it after :) when i am not sleepy ^^